OpenStudy (anonymous):

Find the mean and standard deviation for the data set of the number of elevator stops for several rides. {0, 3, 1, 1, 0, 5, 1, 0, 3, 0}

OpenStudy (anonymous):

The mean is the average.

OpenStudy (anonymous):

mean=14/10=1.4

OpenStudy (anonymous):

standard deviation is square root of variance.

OpenStudy (anonymous):

standard deviation = SR[(x-mean)^2/n]

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

yeah tai u r ryt..:)

OpenStudy (anonymous):

mean=1.4, standard deviation=1.624807681...

OpenStudy (anonymous):

standard deviation = \[\sqrt{\Sigma(x-mu)^2/n}\]